OpenStudy (anonymous):

If the point P(-4/6,y) lies on the unit circle and P is in the third quadrant, what does y equal? If necessary, use the slash mark ( / ) for a fraction bar.

OpenStudy (anonymous):

Oops its P(-4/5,y)

OpenStudy (anonymous):



OpenStudy (anonymous):

Where did you get 20?

OpenStudy (anonymous):

\[\left( \frac{ -4 }{ 5 } \right)^2+y^2=1,y^2=1-\frac{ 16 }{ 25 }=\frac{ 25-16 }{ 25 }=\frac{ 9 }{ 25 }\] \[y=\pm \frac{ 3 }{ 5 }=-\frac{ 3 }{ 5 },\because~y~is~negative. \] In 3rd quadrant y is negative.